# Emergence and Evolution of Interpretable Concepts in Diffusion Models

### v1

@@ 4.4.1 Early-stage interventions
First, we apply spatially targeted interventions according to Eq. 2 using an SAE trained on cond activations of mid_block at $t=1.0$. We observe that a large intervention strength $\beta$ is needed to successfully control the spatial composition consistently. We hypothesize that the skip connections in the U-Net architecture and the features from the null-text conditioning in classifier-free guidance reduce the effect of our interventions, as they provide paths that bypass the intervention. Thus, a larger value of intervention strength is needed to mask the leakage effects. In Eq. 8, we observe that the objects of interest are successfully guided to their respective locations. Moreover, the concepts that we do not intervene on, such as the flower in the first row are preserved.
@@
Figure 8: Effect of spatially targeted interventions at different stages of diffusion, aimed at manipulating image layout. We can efficiently restrict objects to the specified quadrant of the image when intervened in early stages of diffusion. However, in middle and final stages our interventions are unsuccessful.
@@
Next, we perform global interventions according to Eq. (3) aimed at manipulating image style. Interestingly, as depicted in Figure 9, we find that instead of controlling image style, these global interventions broadly modify the composition of the image, without imbuing it with a particular style. As depicted in Figure 10, this phenomenon holds for a wide range of $\beta$. As we vary the intervention strength, we obtain images with various compositions, but without the target style. This observation is consistent with our hypothesis that early stages of diffusion are responsible for shaping the image composition, whereas more abstract and high-level concepts, such as those related to consistent artistic styles emerge later.
@@ 4.4.2 Middle-stage interventions
In an effort to control image style, we perform global interventions in the middle stages. We show results in Figure 9. We find that the these interventions do not alter image composition as in early stages of diffusion. Instead, we observe local edits more aligned with stylistic changes (cartoon look, sandy texture, smooth straight lines, etc.), while the location of objects in the scene are preserved. Contrasting this with early-stage interventions, we hypothesize that the middle stage of diffusion is responsible for the emergence of more high-level and abstract concepts whereas the image layout is already determined in the earlier time steps (also supported by our semantic segmentation experiments). Moreover, varying the intervention strength impacts the intensity of style transfer in the output image (Figure 10).
@@ 4.4.3 Final-stage interventions
Similarly, we find that our global intervention technique is ineffective in manipulating image style in the final stage of diffusion (Figure 9), as we only observe minor textural changes across a wide range of intervention strengths (Figure 10).
(a) Early-stage intervention
(b) Middle-stage intervention
(c) Final-stage intervention
Figure 10: Effect of intervention strength. We perform global intervention on a concept (#$\#1722$ for all time steps) corresponding to cartoon look in top activating images. Early-stage interventions, at any strength, are unable to modify image style consistently but broadly influence image composition. Interventions in the middle stages imbue the image with the target style with increasing intensity. We only observe minor textural changes in final stages of diffusion, even at high intervention strengths.
@@ Appendix A Additional details on SAE training
We train SAEs on the residual updates in the diffusion U-Net blocks down_blocks.2.attentions.1, mid_block.attentions.0, up_blocks.1.attentions.0, referred to as down_block, mid_block, up_block. We use guidance scale of $\omega=7.5$ and $50$ DDIM steps to collect the activations. The dimension of the activation tensor is $16\times 16\times 1280$ for down_block and up_block, $8\times 8\times 1280$ for the mid_block. Training hyperparameters are as follows:
@@
We keep track of normalized mean-squared error (MSE) and explained variance of the SAE reconstructions. In Eq. (1) we provide the complete set of training metrics for all combinations of block, conditioning, timestep, and $k$.
@@ Appendix B Additional details on interventions
In Table 2 we provide the intervention strength ($\beta$) we use for each reverse diffusion stage and intervention type.
@@ Appendix D Qualitative assessment of activations
We visualize the activation maps for top $10$ (in terms of mean activation across the spatial dimensions) activating concepts for generated samples in Figures 12 - 14 for various time steps and blocks. Based on our empirical observations, the activations can be grouped in the following categories:
@@
  • Context-free – We observe that some concepts fire exclusively in specific, structured regions of the image, such as particular corners or bordering edges of the image,irrespective of semantics (see e.g. the last activation in the first row of Figure 12). We hypothesize that these concepts may be a result of optimization artifacts, and are leveraged as semantic-independent knobs for the SAE to reduce reconstruction error. Specifically, if the SAE is unable to "find" $k$ meaningful concepts in the image, as encouraged by the training objective, it may compensate for the missing signal energy in these context-free directions.
@@
Appendix E More examples from the concept dictionary
@@
We depict top $5$ activating concepts, extracted from up_blocks.1.attentions.0, for generated images and their corresponding concept dictionary entries in Figures 15 - 16.
@@
Appendix F Context-free activations
@@
Appendix G Visualization of top dataset examples
@@
Figure 28: Top activating dataset examples for the concept ID 00 belonging to the SAE trained on the cond activation of mid_block at $t=0.0$.

### v2

@@ 4.3.1 Early-stage Interventions
First, we apply spatially targeted interventions according to Eq. 2 using an SAE trained on cond activations of mid_block at $t=1.0$. We observe that a large intervention strength $\beta$ is needed to successfully control the spatial composition consistently. We hypothesize that the skip connections in the U-Net architecture and the features from the null-text conditioning in classifier-free guidance reduce the effect of our interventions, as they provide paths that bypass the intervention. Thus, a larger value of intervention strength is needed to mask the leakage effects. In Figure 4, we observe that the objects of interest are successfully guided to their respective locations. Moreover, the concepts that we do not intervene on, such as the flower in the first row are preserved.
@@
Figure 4: Effect of spatially targeted interventions at different stages of diffusion, aimed at manipulating image layout. We can restrict objects to the specified quadrant of the image when intervened in early stages of diffusion. However, in middle and final stages our interventions are unsuccessful.
@@
Next, we perform global interventions according to Eq. (3) aimed at manipulating image style. Interestingly, as depicted in Figure 5, we find that instead of controlling image style, these global interventions broadly modify the composition of the image, without imbuing it with a particular style. As depicted in Figure 6 (top row), this phenomenon holds for a wide range of $\beta$. As we vary the intervention strength, we obtain images with various compositions, but without the target style. This observation is consistent with our hypothesis that early stages of diffusion are responsible for shaping the image composition, whereas more abstract and high-level concepts, such as those related to consistent artistic styles emerge later.
@@ 4.3.2 Middle-stage Interventions
In an effort to control image style, we perform global interventions in the middle stages. We show results in Figure 5. We find that the these interventions do not alter image composition as in early stages of diffusion. Instead, we observe local edits more aligned with stylistic changes (cartoon look, sandy texture, smooth straight lines, etc.), while the location of objects in the scene are preserved. Contrasting this with early-stage interventions, we hypothesize that the middle stage of diffusion is responsible for the emergence of more high-level and abstract concepts whereas the image layout is already determined in the earlier time steps (also supported by our semantic segmentation experiments). Moreover, varying the intervention strength impacts the intensity of style transfer in the output image (Figure 6 (middle row)).
@@ 4.3.3 Final-stage Interventions
Similarly, we find that our global intervention technique is ineffective in manipulating image style in the final stage of diffusion (Figure 5), as we only observe minor textural changes across a wide range of intervention strengths (Figure 6 (bottom row)).
Figure 6: Effect of intervention strength. We perform global intervention on a concept (#1722\#1722 for all time steps) corresponding to cartoon look in top activating images. Early-stage interventions, at any strength, are unable to modify image style consistently but broadly influence image composition. Interventions in the middle stages imbue the image with the target style with increasing intensity. We only observe minor textural changes in final stages of diffusion, even at high intervention strengths.
@@ 4.4 Summary of observations
While these observations are based on the SDv1.4 diffusion model, we expect similar behavior from different models and architectures. In partiular, we hypothesize that the concepts that are learned by the model are mostly due to the diffusion objective rather than the precise choice of denoising architecture. As noted by Fuest et al. [10], the denoising objective in diffusion models encourages the learning of semantic image representations useful for downstream tasks, much like Denoising Autoencoders (DAEs). The key distinction is that diffusion models condition on the timestep $t$, effectively functioning as a hierarchy of DAEs operating at multiple noise levels. Building on this close analogy between DAEs and diffusion models (rigorously analyzed in [51]), we believe that even if the denoising model architecture changes (such as DiT used in FLUX models vs. the U-Net in SDv1.4, SDXL, etc.), similar underlying concepts should emerge when probed appropriately due to the fundamentally related representations learned by diffusion models.
We also include more quantitative experiments demonstrating that concepts become more refined and distinct as generation progresses (Appendix E.1), and assessing the success of global and spatially targeted interventions at different diffusion stages which conforms with our summary of observations in this section (Appendix E.2).
@@ Appendix A Additional details on SAE training
We obtain the learnable parameters of SAE by optimizing the reconstruction error
where $\alpha$ is a hyperparameter. We use a filtered version of the LAION-COCO [^2] https://huggingface.co/datasets/guangyil/laion-coco-aesthetic, license: apache-2.0 dataset for training prompts. We train SAEs on the residual updates in the diffusion U-Net blocks down_blocks.2.attentions.1, mid_block.attentions.0, up_blocks.1.attentions.0, referred to as down_block, mid_block, up_block. We use guidance scale of $\omega=7.5$ and $50$ DDIM steps to collect the activations. The dimension of the activation tensor is $16\times 16\times 1280$ for down_block and up_block, $8\times 8\times 1280$ for the mid_block. We train all models with Adam optimizer on a single NVIDIA RTX A6000 GPU. Training on $200k$ prompts takes $\approx 1$ hour for up_block and down_block, and $\approx 20$ minutes for mid_block. Training hyperparameters are as follows:
@@
We keep track of normalized mean-squared error (MSE) and explained variance of the SAE reconstructions. In table˜1 we provide the complete set of training metrics for all combinations of block, conditioning, timestep, and $k$.
@@ Appendix B Additional details on interventions
In table˜2 we provide the intervention strength ($\beta$) we use for each reverse diffusion stage and intervention type.
@@
In fig.˜7, we perform global intervention on concept #1722\#1722 (corresponding to cartoon look in top activating images) for all timesteps for various intervention strength $\beta$.
(a) Early-stage intervention
(b) Middle-stage intervention
(c) Final-stage intervention
Figure 7: Effect of intervention strength. We perform global intervention on a concept (#1722\#1722 for all time steps) corresponding to cartoon look in top activating images. Early-stage interventions, at any strength, are unable to modify image style consistently but broadly influence image composition. Interventions in the middle stages imbue the image with the target style with increasing intensity. We only observe minor textural changes in final stages of diffusion, even at high intervention strengths.
@@ Appendix D Qualitative assessment of activations
We visualize the activation maps for top $10$ (in terms of mean activation across the spatial dimensions) activating concepts for generated samples in Figures 9 - 11 for various time steps and blocks. Based on our empirical observations, the activations can be grouped in the following categories:
@@
  • Context-free – We observe that some concepts fire exclusively in specific, structured regions of the image, such as particular corners or bordering edges of the image, irrespective of semantics (see e.g. the last activation in the first row of Figure 9). We hypothesize that these concepts may be a result of optimization artifacts, and are leveraged as semantic-independent knobs for the SAE to reduce reconstruction error. Specifically, if the SAE is unable to "find" $k$ meaningful concepts in the image, as encouraged by the training objective, it may compensate for the missing signal energy in these context-free directions. Visual examples and further discussion can be found in Appendix G.
@@
Appendix E Quantitative results
E.1 Quantifying temporal evolution of concepts
In order to understand how much each SAE feature aligns with a particular concept beyond the qualitative results we have, we introduce two metrics: concept cohesion and concept separability. We refer to concept cohesion as a measure of how tightly clustered the items in the concept dictionary are around their concept center. Similarly, we calculate concept separability as a measure of how distinct concepts are from one another. Next, we talk about in detail how we calculate these scores. To calculate concept cohesion, for each concept, we calculate the mean word2vec embedding of the items in the concept dictionary. Then, for each item in the dictionary entry, we calculate the cosine distance to the mean embedding and take the average. The final score is calculated by taking the average across concepts. As for the concept separability, we first calculate the concept centers as described before. Note that, we have an $n\times d$ tensor where $n$ is the number of concepts in our dictionary and $d$ is the word2vec embedding dimension. We then calculate the pairwise cosine similarity matrix (of shape $n\times n$). Then we denote the final score as the average of non-diagonal entries. We calculate these for our concept dictionaries for each timestep $t\in\left\{0.0,0.5,1.0\right\}$. The results are presented in Table 3.
Table 3: Concept cohesion and separability across timesteps $t$.
$t$ Concept Cohesion Concept Separability
0.0 $0.664\pm 0.182$ $0.344\pm 0.159$
0.5 $0.639\pm 0.170$ $0.363\pm 0.154$
1.0 $0.588\pm 0.169$ $0.433\pm 0.169$
We observe that as we move along the reverse diffusion iterates ($t=1.0$ to $t=0.0$), the concepts that we have identified become purer (as measured by the increase in concept cohesion from $0.588$ to $0.664$) Moreover, average pair-wise similarity of concepts decrease from $0.433$ to $0.344$. This indicates that concepts are becoming more distinct and separated as generation progresses.
E.2 Quantitative assessment of intervention success
Spatially targeted interventions – To quantitatively assess the success of spatially targeted interventions summarized in Figure 4, we utilize CLIPSeg [lüddecke2022imagesegmentationusingtext] as a zero-shot segmentation method to determine whether the object in the generated image is in the intended quadrant. More specifically, using the object name, CLIPSeg gives a prediction score for each pixel in the generated image. We calculate the center of the mass and determine which quadrant of the image it belongs to. We run the experiments on an extended object list consisting of 10 in total ("bee", "book", "dog", "apple", "banana", "cat", "car", "phone", "door", "bird"). We select these objects as they are commonly observed in the LAION-COCO dataset. For each timestep, we calculate the overall spatial edit success as the average matching score across all objects and $4$ quadrants. We present the average score in Table 4.
Table 4: Performance of spatially targeted interventions across timesteps $t$.
Method $t=0.0$ $t=0.5$ $t=1.0$
Random Baseline 0.25 0.25 0.25
Ours 0.23 0.35 0.80
We observe the edits at timesteps $t=0.0$ and $t=0.5$ unsuccessful and to be comparable to the random prediction baseline. At $t=1.0$, our edits are successful $80\%$ of the time. It is likely that we can improve the performance further with hyperparameter tuning, which we leave it for future work.
Global interventions – Since no ground truth images are available in the case of global interventions, we adopt a proxy metric to measure success (inspired by earlier editing-focused works such as Prompt-to-Prompt [15]): CLIP similarity between the generated image and the mean CLIP embedding of the target concept (computed as the average embedding of all dictionary words associated with that concept). We measure this similarity both before and after the intervention. An edit is considered successful if the similarity increases post-intervention. We also report the average LPIPS score between the image before and after intervention as a proxy to the quality of the generated image. That is, ideally we want the edited images to stay close to the “reference” (before intervention) image.
To perform this analysis, we sample $1000$ prompts from the validation set for added robustness and apply our editing method at three different diffusion timesteps (early, middle, late) for the concept indices highlighted in Figure 5 (which correspond to the concepts cartoon, sea & sand, and painting). The results are summarized in Table 5.
Table 5: Global intervention performance across timesteps $t$.
[…]
$t=0.0$ 1722 0.201 0.208 +0.007 0.85 0.114
2349 0.190 0.194 +0.004 0.73 0.094
3593 0.201 0.206 +0.005 0.76 0.133
Average 0.197 0.203 +0.006 0.78 0.114
$t=0.5$ 1722 0.203 0.222 +0.019 0.94 0.379
524 0.192 0.220 +0.028 0.96 0.407
2137 0.203 0.218 +0.015 0.88 0.369
Average 0.199 0.220 +0.021 0.93 0.385
$t=1.0$ 1722 0.202 0.212 +0.010 0.77 0.647
2366 0.194 0.197 +0.003 0.58 0.646
$29$ 0.203 0.212 +0.009 0.73 0.665
Average 0.200 0.207 +0.007 0.69 0.653
[…]
Appendix F Additional details on building concept dictionary and predicting image composition
F.1 Extracting interpretations from SAE features
Figure 12: Curating the concept dictionary: 1) We cache SAE activations across time steps and blocks during generation. 2) We leverage a pipeline of image tagging, open-set object detection and promptable segmentation to annotate the generated image with segmentation masks and object labels. 3) We find SAE activations that overlap with the object masks. 4) We add the overlapping object’s label to the concept dictionary under the matching SAE activation’s CID.
In Figure 12, we provide a detailed figure to depict the curation of the concept dictionary (also explained in Section 3.3). To build the concept dictionary, we first sample a set of text prompts, generate the corresponding images using a diffusion model and extract the SAE activations for each CID during generation. We obtain ground truth annotations for each generated image using a pre-trained vision pipeline, that combines image tagging, object detection and semantic segmentation, resulting in a mask and label for each object in generated images. Finally, we evaluate the alignment between our ground truth masks and the SAE activations for each CID, and assign the corresponding label to the CID only if there is sufficient overlap.
F.2 Predicting image composition from SAE features
In Figure 13, we show the pipeline that we use to predict image composition in detail. Leveraging the concept dictionary, we predict the final image composition based on SAE features at any time step, allowing us to gain invaluable insight into the evolution of image representations in diffusion models. Suppose that we would like to predict the location of a particular object in the final generated image, but before the reverse diffusion process is completed. First, given SAE features from a given intermediate time step, we extract the top activating concepts for each spatial location. Next, we create a conceptual map of the image by assigning a word embedding to each spatial location based on our curated concept dictionary. This conceptual map shows how image semantics, described by localized word embeddings, vary spatially across the image. Given a concept we would like to localize, such as an object from the input prompt, we produce a target word embedding and compare its similarity to each spatial location in the conceptual map. To produce a predicted segmentation map, we assign the target concept to spatial locations with high similarity, based on a pre-defined threshold value. This technique can be applied to each object present in the input prompt (or to any concepts of interest) to predict the composition of the final generated image.
Figure 13: Predicting image composition: 1) We cache SAE activations at the first diffusion step (or other time step of interest) and extract top activated concepts per spatial location. 2) We fetch corresponding objects from the concept dictionary and produce a conceptual embedding via Word2Vec. 3) We compare the conceptual embedding at each location to the target word embeddings from the input prompt and predict a segmentation map based on cosine similarity.
F.3 More examples from the concept dictionary
@@
We depict top $5$ activating concepts, extracted from up_blocks.1.attentions.0, for generated images and their corresponding concept dictionary entries in Figures 14(b) - 16.
(a) $t=1.0$
(b) $t=0.0$
Figure 14: Concept dictionary and visualization of the activation maps for the top $5$ activating concepts. Sample ID: $2000031$
@@
Appendix G Context-free activations
@@
Appendix H Visualization of top dataset examples
@@
Figure 28: Top activating dataset examples for the concept ID 0 belonging to the SAE trained on the cond activation of mid_block at $t=0.0$.
@@
Appendix I Broader Impact Statement
In this work, we demystify the inner workings of text-to-image diffusion models by revealing how visual representations evolve over the reverse diffusion process. By uncovering interpretable features and developing stage-specific editing techniques, our approach may enable more granular and controllable image generation, which can benefit safety critical applications. However, we acknowledge that greater interpretability and control also carry risks, such as facilitating misuse for deceptive content generation or circumventing safety mechanisms.